OpenStudy (anonymous):

Tiles have weights that are independently normally distributed with a mean of 45.3 and a standard deviation of 0.02. What is the probability that the total weight of three tiles is no more than 135.975.

ganeshie8 (ganeshie8):

so basically we need tofind : \(P(x_1 + x_2 + x_3 \le 135.975)\)

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

I tried fininding just the probability of 1 tile but that didn't work. I would multiply that probability by 3 but that wont work.

ganeshie8 (ganeshie8):

average value will do

ganeshie8 (ganeshie8):

finding \(P(x_1 + x_2 + x_3 \le 135.975)\) is same as finding below : \(\large P(\overline{x} \le \frac{135.975}{3}) \)

OpenStudy (anonymous):

That's what I did!!

ganeshie8 (ganeshie8):

\(\overline{x}\) is the average weight

OpenStudy (anonymous):

But my probability i over 100% :( .

ganeshie8 (ganeshie8):

lets see..

OpenStudy (anonymous):